$SPAD/src/input fr2.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool fr2.output
)set message test on
)set message auto off
)clear all
--S 1 of 6
double(x) == x + x
--R
--R                                                                   Type: Void
--E 1
--S 2 of 6
f := factor(720)
--R
--R
--R         4 2
--R   (2)  2 3 5
--R                                                      Type: Factored(Integer)
--E 2
--S 3 of 6
map(double,f)
--R
--R   Compiling function double with type Integer -> Integer
--R
--R           4 2
--R   (3)  2 4 6 10
--R                                                      Type: Factored(Integer)
--E 3
--S 4 of 6
makePoly(b) == x + b
--R
--R                                                                   Type: Void
--E 4
--S 5 of 6
g := map(makePoly,f)
--R

--R   Compiling function makePoly with type Integer -> Polynomial(Integer)
--R
--R
--R                      4       2
--R   (5)  (x + 1)(x + 2) (x + 3) (x + 5)
--R                                          Type: Factored(Polynomial(Integer))
--E 5
--S 6 of 6
nthFlag(g,1)
--R
--R
--R   (6)  "nil"
--R                                                       Type: Union("nil",...)
--E 6
)spool
)lisp (bye)

———-

References^{[1] nothing}

5